in a vector space v over R, prove that (-1)X = -7 Hind (you may use the fact that Ox=0)
Answer #1

For any queries please comment below otherise, give a thumb's up.Answer :- let v be a vector space over TR To shoul (-1) = -X Noul (l)x+(1) X [-1) +(1)] = [-1+1)x 0.x = 0 as 0.X 10+0) X = 0.F